OpenStudy (anonymous):

A pebble is tossed from the top of a cliff. The pebble's height in feet is given by y(t) = -16t2 + 6t + 200, where t is the time in seconds. Its horizontal distance in feet from the base of the cliff is given by d(t) = 5t. How far will the pebble be from the base of the cliff when it hits the ground? Round to the nearest foot. _______ ft Use a number only.

OpenStudy (campbell_st):

the vertical distance will be zero so y(t) = 0 16t^2 -6t - 200 = 0 2(8t^2 - 3t - 100) = 0 using the general quadratic formula \[t = (3 \pm \sqrt{3209})/16\] only consider the positive t value \[t = (3 + \sqrt{3209})/16\] Substitute into the horizontal distance equation \[d(t) = 5 \times (3 \pm \sqrt{3209})/16\] the pebble will be 19 ft from the base of the cliff

OpenStudy (anonymous):

oh i get it thx